# Dynamics of states of infinite quantum systems as a cornerstone of the second law of thermodynamics

### v2

@@ 2.4 Dynamical properties of states of infinite systems related to the modified second law
is a mathematically precise definition of a sudden interaction: it represents a “finite discontinuity” which replaces the (mathematically) ill-defined $\delta(t)$- infinity” of the lifting of barriers in constrained equilibria ([Pa33]), such as the barrier model of remark  2.1.
@@ 2.4.3 Quantum chaotic states: the role of dynamics and the classical character
denote the generalized ferromagnetic Ising model (gIm), and $\tau$ the corresponding automorphism corresponding to the infinite system. Then, for both the exponential model  (25), and the Dyson model  (26) with $\alpha>2$, there exists an infinite number of distinct adiabatic transformations satisfying  (53), i.e., associated to sudden perturbations, each of extended type, such that modified second law (Clausius) for $T=0$ holds. For perturbations of local type of the ferromagnetic anisotropic Heisenberg model  (23), the modified second law (Clausius) does not hold.
@@
That is to say,  (53) is satisfied, characterizing a sudden interaction, where $\tau_{1}$ denotes the automorphism defined by
$\tau_{1}\equiv\tau\circ\tau_{P}=\tau_{P}\circ\tau$ (71)
and $\tau_{P}$ is the automorphism associated to a perturbation of the original system, associated to a finite-volume operator $V_{\Lambda}$ which is assumed to commute with the finite-region Hamiltonian  (20), i.e.,
$V_{\Lambda},H_{\Lambda}]=0\forall\Lambda\subset\mathbf{Z}^{\nu}$ (72)
Above, then, $\tau_{P}$ is defined by
$\tau_{P,t}(A)\equiv\lim_{\Lambda\nearrow\mathbf{Z}^{\nu}}(\exp(itV_{\Lambda})A\exp(-itV_{\Lambda}))\forall A\in{\cal A}_{\Lambda}$ (73)
and then by extension as usual. Take $\omega=\omega_{0}^{+}$ as one of the ground states of the Hamiltonian defined by  (69), the other case, $\omega_{0}^{-}$, as well as any convex combination of the two, being similar. Choose, in  (73), $V_{\Lambda}$ as a tranverse, sufficiently strong magnetic field (which commutes with $H_{\Lambda}$):
@@
If $A$ and $B$ are chosen as positive integers, by  (79) $f$ ranges over all rational numbers between zero and one. With our choice $T=0$, the restriction of the initial ground state to a region $\Lambda$, $\omega_{\Lambda,0}^{+}$, shifts under conditions  (74) to  (80), to the ground state of $V_{\Lambda}$, which is the vector state $(\Psi_{\Lambda}^{f},\cdot\Psi_{\Lambda}^{f})$, with
@@
Above, by definition, $\sigma_{x}^{1}|\pm)=\pm\frac{1}{2}|\pm)$, and the sets $A_{k}$ and $B_{k}$ have been defined in  (77),  (78) and  (79), and are thus seen to satisfy $A_{k}\cup B_{k}=[kC+1,(k+1)C]$. Moreover, by  (79) and  (80), and the fact that $A$ and $B$ are positive integers, it follows that $C>1$. The limiting state

### v3

@@ 2.4 Dynamical properties of states of infinite systems related to the modified second law
is a mathematically precise definition of a sudden interaction: it represents a “finite discontinuity” which replaces the $\delta(t)$- infinity” of the lifting of barriers in constrained equilibria ([Pa33]), such as the barrier model of remark  2.1. In the examples, however, the infinity appears as a “coupling constant” associated to the perturbation.
@@ 2.4.3 Quantum chaotic states: the role of dynamics and the classical character
denote the generalized ferromagnetic Ising model (gIm), and $\tau$ the corresponding automorphism corresponding to the infinite system. Then, for both the exponential model  (25), and the Dyson model  (26) with $\alpha>2$, there exists an infinite number of distinct adiabatic transformations satisfying  (53), i.e., associated to a class of sudden perturbations, each of extended type, such that modified second law (Clausius) for $T=0$ holds. For perturbations of local type of the ferromagnetic anisotropic Heisenberg model  (23), the modified second law (Clausius) does not hold.
@@
That is to say,  (53) is satisfied, characterizing a sudden interaction, where $\tau_{1}$ denotes the automorphism defined by a finite-volume Hamiltonian
$H(\lambda,\Lambda)\equiv H_{\Lambda}+\lambda V_{\Lambda}$ (71)
Above, $V_{\Lambda}$ is a perturbation, and $\lambda$ a coupling constant. We take $\omega=\omega_{0}^{+}$ as one of the ground states of the Hamiltonian defined by  (69), the other case, $\omega_{0}^{-}$, as well as any convex combination of the two, being similar. Choose the perturbation $V_{\Lambda}$ as a tranverse, sufficiently strong magnetic field
@@
If $A$ and $B$ are chosen as positive integers, $f$ ranges over all rational numbers between zero and one.
With our choice $T=0$, the restriction of the initial ground state to a region $\Lambda$, $\omega_{\Lambda,0}^{+}$, shifts under conditions above ( (71) et seq.), in the limit of infinite coupling
$\lambda\to\infty$ (79)
to the ground state of $V_{\Lambda}$, which is the vector state $(\Psi_{\Lambda}^{f},\cdot\Psi_{\Lambda}^{f})$, with
@@
In fact, the limit of infinite coupling is precisely defined by the above conditions.
Above, by definition, $\sigma_{x}^{1}|\pm)=\pm\frac{1}{2}|\pm)$, and the sets $A_{k}$ and $B_{k}$ satisfy $A_{k}\cup B_{k}=[kC+1,(k+1)C]$. Moreover, it follows that $C>1$. The limiting state